OpenStudy (phebe):

FANSSS+MEDALSSS!!! Solve. b. √(2x+4) = √3x

OpenStudy (phebe):

@undeadknight26

OpenStudy (solomonzelman):

raise both sides to the second power. What will you get after doing that?

OpenStudy (phebe):

which both sides

OpenStudy (solomonzelman):

raise each side to the second power. (in this case means to remove the square root symbols)

OpenStudy (phebe):

ok

OpenStudy (solomonzelman):

what do you get after doing that ?

OpenStudy (phebe):

2^2=4 and 3^2=9

OpenStudy (solomonzelman):

2x+4=3x is what you get, if you just remove the √ from each side.

OpenStudy (solomonzelman):

Then subtract 2x from each side.

OpenStudy (phebe):

ok

OpenStudy (phebe):

2x-2 cancelled out and 3-2x=1

OpenStudy (phebe):

@SolomonZelman

OpenStudy (phebe):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{purple }{ \rm 2x+4= 3x}\) \(\large\color{purple }{ \rm 4= x}\)

OpenStudy (phebe):

is 4 tyhe answer @SolomonZelman